# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ 3 Splitting & convergence speed: multivariate Gaussian case
    Assume the conditions of (i) or (ii). Then the asymptotic (large $n$) ratio of the $\mathrm{KL}$ distance to $\pi$ of the W-FR split scheme solution with respect the solution of the exact WFR is given by
    $\lim_{n\rightarrow\infty}\frac{\mathrm{KL}(\nu_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}=\frac{p_{1}^{\top}(E_{0}^{-1}C_{\pi}+\Omega^{\beta})^{-1}D_{0}(E_{0}^{-1}C_{\pi}+\Omega^{\beta})^{-1}p_{1}}{p_{1}^{\top}(E_{0}^{-1}C_{\pi}+\Omega)^{-1}D_{0}(E_{0}^{-1}C_{\pi}+\Omega)^{-1}p_{1}},$ (3.18)
    where $\Omega$ and $\Omega^{\beta}$ are as defined in (3.11) and (3.14) respectively, $p_{1}$ is the eigenvector corresponding to the smallest eigenvalue of $\Gamma$ and $D_{0}:=E_{0}^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}E_{0}^{-1}C_{\pi}$. Likewise, for the FR-W split scheme, the same limiting ratio for $\frac{\mathrm{KL}(\eta_{n\gamma}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}$ holds with $\Omega^{\beta}$ replaced by $\Omega^{\alpha}$ as defined in (3.17).
Notice that in dimension 1, the limiting ratio in (3.18) is independent of $\varepsilon_{0}$, so that the speed-up depends only on the initial covariance relative to the target covariance. In particular, in the 1-dimensional case if $C_{0}>C_{\pi}$ then FR-W will converge faster than the exact WFR dynamics, conversely, if $C_{\pi}>C_{0}$ then W-FR will be faster. In the multivariate case however, the squared mean difference $\varepsilon\varepsilon^{\top}$ plays an important role also.
@@
Figure 3.2: Left: Ratio of KL from $n$-step W-FR scheme to KL from exact WFR as a function of $t=n\times\gamma$, $\pi$ is a 10D Gaussian (see Appendix B [us_neurips] for details) and $\gamma=0.7$. The horizontal black line corresponds to (3.18). For reference, $\mathrm{KL}(\mu_{4.2}||\pi)=9.7$. Right: Same as left but for FR-W scheme.
4 Preservation of log-concavity of exact WFR
The previous section shows that, in the Gaussian case, splitting schemes can speed up convergence provided $\mu_{0}$ is chosen wisely.
In order to generalise this result to say, log-concave target distributions, we first provide a result demonstrating that log-concavity is preserved under the evolution of the exact WFR flow (2.1). Recall that for any log-concave $\mu_{0}$, preservation of log-concavity uniformly in time under the W flow is guaranteed only for $\pi$ Gaussian (i.e. for Ornstein-Uhlenbeck Semigroups) due to [Kolesnikov2001]. This is in stark contrast to the FR flow, where log-concavity is preserved uniformly in time for any strongly log-concave $\pi$ and $\mu_{0}$ (see Lemma C.1).
With some stronger conditions on $\mu_{0}$ and $\pi$, as detailed in Assumptions 2 and 3, we establish preservation of log-concavity under the W flow for a limited time horizon, using similar arguments as in [Pathiraja2021, Section 7.2] (in the context of stochastic filtering) and Lemma 16 in [liang_characterizing_2025]. We first state the necessary assumptions:
@@ 4 Preservation of log-concavity of exact WFR
    Denote by $R:=-\Delta V_{\pi}+|\nabla V_{\pi}|^{2}$. Also define $\mathcal{H}:=R+V_{\pi}$. Assume $\mathcal{H}$ is strongly convex with parameter $\alpha_{h}>0$. This implies that
    $\nabla^{2}R\succeq(\alpha_{h}-L_{\pi})I,$
    where $\alpha_{h}-L_{\pi}<0$.
@@
W flows preserve log-concavity over a fixed time horizon. Suppose $\mu_{t}$ is the solution of the W flow (2.2) at time $t$, initialised at $\mu_{0}(x)$ satisfying Assumption 3. Let $b:=\sqrt{|\alpha_{h}-L_{\pi}|}$. Then there exists a $t^{\ast}>0$ such that for all $t<t^{\ast}$, $\mu_{t}(x)\propto e^{-\mathcal{E}_{t}(x)}$, with $\mathcal{E}_{t}(x)$ a strongly convex function,
$\nabla^{2}\mathcal{E}_{t}(x)$ $\succeq\left(\frac{\alpha_{\pi}}{2}+c_{t}\right)I,\quad\forall\enskip x\in\mathbb{R}^{d},$ (4.1)
$c_{t}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-bt\right)>0,$ (4.2)
$c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi}.$ (4.3)
Firstly, notice that in Lemma 4.1, we consider the case where $\alpha_{h}-L_{\pi}<0$, since $\alpha_{h}>L_{\pi}$ implies $R$ is strongly convex. Strong convexity of $R$ would imply uniform in time convexity of $\mu_{t}$ by similar techniques as in the proof of Lemma 4.1, which is not guaranteed unless $\pi$ is Gaussian due to [Kolesnikov2001]. The length of the time horizon over which strong convexity is preserved also depends on the log-concavity of the difference between initial and target potentials.
In the next theorem, we show that the preservation of log-concavity under the FR flow can be exploited to ensure the same for the WFR flow, uniformly in time. The proof of the theorem can be found in Appendix C.2.
@@
WFR preserves log-concavity uniformly in time. Assume the conditions of Lemma 4.1 and that $b:=\sqrt{\frac{|\alpha_{h}-L_{\pi}|}{2}}$ satisfies
$b^{2}<\tfrac{\alpha_{\pi}}{2}.$ (4.4)
@@
$\alpha_{t}=\frac{\alpha_{\pi}}{2}+\frac{c_{\infty,1}-c_{\infty,2}l_{0}e^{-t(c_{\infty,1}-c_{\infty,2})}}{1-l_{0}e^{-t(c_{\infty,1}-c_{\infty,2})}},$ (4.5)
and
$l_{0}$ $=\frac{(c_{0}-c_{\infty,1})}{(c_{0}-c_{\infty,2})};\qquad c_{\infty,1}=\frac{-1+\sqrt{1+4(\frac{\alpha_{\pi}}{2}-b^{2})}}{2};\qquad c_{\infty,2}=\frac{-1-\sqrt{1+4(\frac{\alpha_{\pi}}{2}-b^{2})}}{2};$
and $c_{0}$ is given by (4.3).
The tightness of the constants (4.5) is evaluated empirically for the 1D Gaussian case in Figure 4.1 where the true constants are known. As can be seen there, for a sufficiently large difference in convexity of the initial and target potentials, the constants (4.5) are reasonably tight (c.f. left vs right). We will rely on the results of Theorem 4.1 in the next section to characterise $\mathrm{KL}$ convergence rates.
Figure 4.1: Comparison of the log-concavity constant predicted by Theorem 4.1 and the true log-concavity constant obtained from (3.1) for a 1D Gaussian with $\mu_{0}(x)=\mathcal{N}(x;,0,1)$ and $\pi(x)=\mathcal{N}(x;,0,100)$ (left), $\pi(x)=\mathcal{N}(x;,0,5)$ (middle), $\pi(x)=\mathcal{N}(x;,0,2.1)$ (right). For $C_{\pi}\leq 2$ the assumptions of Theorem 4.1 are not satisfied.
5 Convergence speed of Exact WFR: log-concave case
The rate of decay of $\mathrm{KL}$ for the WFR flow is obtained considering the time derivative of $\mathrm{KL}(\mu_{t}||\pi)$ along the WFR flow
@@ 5 Convergence speed of Exact WFR: log-concave case
As observed in [gallouet2017jko, page 13], the first term corresponds to the negative gradient of $\mathrm{KL}$ w.r.t. the Wasserstein-2 metric, while the remaining terms give the negative gradient of $\mathrm{KL}$ w.r.t. the Fisher–Rao geometry, implying that the dissipation for the WFR metric is the sum of the W and the FR dissipation. As a consequence, the convergence rate of WFR gradient flow (2.1) can be upper bounded by
$\mathrm{KL}(\mu_{t}||\pi)\leq\min\left\{\mathrm{KL}(\mu_{t}^{\textrm{FR}}|\pi),\mathrm{KL}(\mu_{t}^{\textrm{W}}|\pi)\right\},$ (5.2)
where $\mu_{t}^{\textrm{FR}},\mu_{t}^{\textrm{W}}$ denote the solutions of the FR and W flow PDEs respectively. A bound on the decay of KL along the W flow can be obtained if $\pi$ satifies an LSI (2.8) (e.g., [chewi2024analysis]). Decay of KL along the FR flow at rate $\mathcal{O}(e^{-t})$ is guaranteed under boundedness conditions on the score $\log\pi/\mu_{0}$ ([chen2023sampling, Theorem 4.1] and [lu2023birth, Proposition 2.4]).
A sharper rate of converge for WFR can be obtained assuming $\log\nicefrac{{\pi(x)}}{{\mu_{0}(x)}}\geq M$ and a warm-start condition [Lu2019, Appendix B]
@@
for all $t\geq t_{0}:=\log(M/\delta^{3})$ and $\delta>0$.
We point out that neither (5.2) nor (5.3) provide a sharp rate of decay for $\mathrm{KL}(\mu_{t}||\pi)$. In fact, (5.2) shows that the rate of WFR is never worse than that of W or FR, looking at (5.1) however it is evident that the rate of decay of the WFR flow should be considerably higher than that of the two singular flows. Similarly, (5.3) shows that the decay in the case of the WFR flow is faster than that of the W flow but the rate depends on an arbitrary constant $\delta$ which makes (5.3) harder to interpret.
In Proposition 5.1, we obtain a result that shows the decay rate is indeed the sum of the W and FR rates, but for the symmetrised KL, not KL. The symmetrised $\mathrm{KL}$ divergence, or Jeffrey’s divergence $J$, is defined as
@@
for any $\mu\ll\pi$ and also $\pi\ll\mu$ . The rationale behind considering the symmetrised KL is due to the fact that the KL is not geodesically convex under the FR flow [carrillo_fisher-rao_2024, Theorem 1.1] nor satisfies a gradient dominance condition [carrillo_fisher-rao_2024, Theorem 4.1], which makes its convergence analysis more difficult. On the other hand, Jeffrey’s divergence satisfies a gradient dominance condition [carrillo_fisher-rao_2024, Section 4] which allows to achieve improved rates. In addition, $J(\mu,\pi)$ upper bounds $\mathrm{KL}(\mu||\pi)$ and thus the obtained rates for $J(\mu,\pi)$ gives information on the decay of KL too. The Kl decay rate obtained in Proposition 5.1 is a sum of the decay rates of W and FR flow, conjectured by [domingo-enrich2023an] but not yet proven in the literature.
Theorem 4.1 immediately provides sufficient conditions to establish an exponential decay result of the symmetrised $\mathrm{KL}$ divergence, or Jeffrey’s divergence $J$. The proof makes use of recent functional inequality results on pure FR flows [carrillo_fisher-rao_2024, Section 4] and provides a sharper rate of convergence than the results available in the literature (Figure 5.1).
@@
Let Assumption 2 hold. Then the following decay result holds for $\mu_{t}$ the solution of the WFR PDE (2.1) for all $t>0$
$J(\mu_{t},\pi)\leq J(\mu_{0},\pi)e^{-t\kappa},$
where $\kappa=2\min(\alpha_{\pi},\alpha_{t})+1$ and $\alpha_{t}$ is given in Theorem 4.1.
@@
Assuming valid interchange of differentiation and integration,
@@
$\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)f_{\textrm{W}}(\mu_{t})dx$ $=-\int\left(\log\frac{\mu_{t}(x)}{\pi(x)}-\frac{\pi(x)}{\mu_{t}(x)}\right)\nabla\cdot\left(\mu_{t}(x)\nabla\log\frac{\mu_{t}}{\pi}(x)\right)(x)dx$ (5.5)
@@
$\leq-2\min(\lambda_{\pi}^{-1},\lambda_{\mu_{t}}^{-1})(\mathrm{KL}(\mu_{t}||\pi)+\mathrm{KL}(\pi||\mu_{t}))$
@@
where $\mathcal{I}(\mu_{t}||\pi)$ is the relative Fisher information between $\mu_{t}$ and $\pi$. The last inequality holds due to Assumption 2 and Theorem 4.1. More specifically, since $\pi$ is $\alpha_{\pi}$-strongly log-concave, it immediately satisfies a Log-Sobolev inequality (2.8) with constant $\lambda_{\pi}=\alpha_{\pi}^{-1}$ (likewise for $\mu_{t}$, which is $\alpha_{t}$-strongly log-concave due to Theorem 4.1). Then for the second term in (5.4), by a direct application of (5.16) in [carrillo_fisher-rao_2024, Theorem 5.6] with $f(y)=y\log(y)$ and $\bar{f}=yf(y^{-1})$, it holds that
$-\int\frac{\pi}{\mu_{t}}f_{\textrm{FR}}(\mu_{t})dx\leq-J(\mu_{t},\pi)l,$
@@
$\int\log\frac{\mu_{t}}{\pi}f_{\textrm{FR}}(\mu_{t})dx=-\text{Var}_{\mu_{t}}\left(\log\frac{\mu_{t}}{\pi}\right)<0,$
@@
$\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)f_{\textrm{FR}}(\mu_{t})dx\leq-J(\mu_{t},\pi).$ (5.6)

### v2

@@ 3 Splitting & convergence speed: multivariate Gaussian case
    Assume the conditions of (i) or (ii) and that $m_{0}\neq m_{\pi}$. Then the asymptotic (large $n$) ratio of the $\mathrm{KL}$ distance to $\pi$ of the W-FR split scheme solution with respect the solution of the exact WFR is given by
    $\lim_{n\rightarrow\infty}\frac{\mathrm{KL}(\nu_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}=\frac{p_{1}^{\top}(I+E_{0}\Omega^{\beta}C_{\pi}^{-1})^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}(I+E_{0}\Omega^{\beta}C_{\pi}^{-1})^{-1}p_{1}}{p_{1}^{\top}(I+E_{0}\Omega C_{\pi}^{-1})^{-1}\varepsilon_{0}\varepsilon_{0}^{\top}(I+E_{0}\Omega C_{\pi}^{-1})^{-1}p_{1}},$ (3.12)
    where $\Omega$ and $\Omega^{\beta}$ are as defined in (3.9) and (3.10) respectively, $p_{1}$ is the eigenvector corresponding to the smallest eigenvalue of $C_{\pi}^{-1}$. Likewise, for the FR-W split scheme, the same limiting ratio for $\frac{\mathrm{KL}(\eta_{n}||\pi)}{\mathrm{KL}(\mu_{n\gamma}||\pi)}$ holds with $\Omega^{\beta}$ replaced by $\Omega^{\alpha}$ as defined in (3.11).
Notice that in dimension 1, the limiting ratio in (3.12) is independent of $\varepsilon_{0}$, and the speed-up depends only on the initial covariance relative to the target covariance. In particular, if $C_{0}>C_{\pi}$ then FR-W will converge faster than the exact WFR dynamics, conversely, if $C_{\pi}>C_{0}$ then W-FR will be faster. In the multivariate case however, the mean difference $\varepsilon_{0}$ plays an important role also and the sign of $E_{0}$ alone cannot be used to determine where the ratio in (3.12) is strictly larger or smaller than 1. Instead, for a chosen step size $\gamma$, if
$|p_{1}^{\top}(I+E_{0}\Omega^{\beta}C_{\pi}^{-1})^{-1}\varepsilon_{0}|<|p_{1}^{\top}(I+E_{0}\Omega C_{\pi}^{-1})^{-1}\varepsilon_{0}|$
then W-FR will converge faster than the exact dynamics (similarly for FR-W with $\Omega^{\beta}$ replaced by $\Omega^{\alpha}$).
@@
Figure 3.2: Left: Ratio of KL from $n$-step W-FR scheme to KL from exact WFR as a function of $t=n\times\gamma$, $\pi$ is a 10D Gaussian and $\gamma=0.7$. The horizontal black line corresponds to (3.12). For reference, $\mathrm{KL}(\mu_{4.2}||\pi)=9.7$. Right: Same as left but for FR-W scheme.
4 Preservation of log-concavity
The previous section shows that, in the Gaussian case, splitting schemes can speed-up convergence provided $\mu_{0}$ is chosen wisely. In order to generalise this result to say, log-concave target distributions, we first provide a result demonstrating that log-concavity is preserved under the evolution of the exact WFR flow (2.1). Recall that for any log-concave $\mu_{0}$, preservation of log-concavity uniformly in time under the W flow is guaranteed only for $\pi$ Gaussian (i.e. for Ornstein-Uhlenbeck Semigroups) due to [KOL01]. This is in stark contrast to the FR flow, where log-concavity is preserved uniformly in time for any strongly log-concave $\pi$ and $\mu_{0}$ (see Lemma C.1).
With some stronger conditions on $\mu_{0}$ and $\pi$, as detailed in Assumptions 2 and 3, we establish preservation of log-concavity under the W flow for a limited time horizon, using similar arguments as in [PRS21, Section 7.2] (in the context of stochastic filtering) and Lemma 16 in [LMW25]. We first state the necessary assumptions:
@@ 4 Preservation of log-concavity
    Denote by $R:=-\frac{1}{2}\Delta V_{\pi}+\frac{1}{4}|\nabla V_{\pi}|^{2}$. Also define $\mathcal{H}:=R+V_{\pi}$. Assume $\mathcal{H}$ is strongly convex with parameter $\alpha_{h}>0$.
@@
W flows preserve log-concavity. Suppose $\mu_{t}$ is the solution of the W flow (2.2) at time $t$, initialised at $\mu_{0}(x)$ satisfying Assumption 3. Let $b:=\sqrt{\frac{|\alpha_{h}-L_{\pi}|}{2}}$. Then for some $0<\delta<1$ as in Assumption 3, the following holds
  1. (i)
    Fixed time horizon: Suppose $\alpha_{h}-L_{\pi}<0$. Then there exists a $t^{\ast}>0$ such that for all $t<t^{\ast}$, $\mu_{t}(x)\propto e^{-\mathcal{E}_{t}(x)}$, with $\mathcal{E}_{t}(x)$ a strongly convex function where $\nabla^{2}\mathcal{E}_{t}(x)\succeq\left(\frac{\alpha_{\pi}}{2}+c_{t}\right)I,\quad\forall\enskip x\in\mathbb{R}^{d},$ where
    $c_{t}$ $=b\tan\left(\tan^{-1}\left(\frac{c_{0}}{b}\right)-2bt\right)$ (4.1)
    $c_{0}$ $=\alpha_{d}+\frac{\delta}{2}\alpha_{\pi},$ (4.2)
  2. (ii)
    Uniform in time: Suppose $\alpha_{h}-L_{\pi}>0$. Then for all $t\geq 0$, $\mu_{t}(x)\propto e^{-\mathcal{E}_{t}(x)}$, with $\mathcal{E}_{t}(x)$ a strongly convex function where $\nabla^{2}\mathcal{E}_{t}(x)\succeq\left(\frac{\alpha_{\pi}}{2}+c_{t}\right)I,\quad\forall\enskip x\in\mathbb{R}^{d},$ where
    $c_{t}=b\left(\frac{K-e^{-4bt}}{K+e^{-4bt}}\right)$ (4.3)
    with $K:=\frac{b+c_{0}}{b-c_{0}}$ and $c_{0}$ as in (4.2).
The case of Gaussian $\mu_{0}$ and $\pi$ is covered by (ii) in the above lemma, where it can be easily verified that $\alpha_{h}-L_{\pi}>0$. More generally, when $\alpha_{h}-L_{\pi}<0$, a uniform in time preservation of log-concavity is not guaranteed, and the length of the time horizon over which strong convexity is preserved depends on the log-concavity of the difference between initial potential $V_{0}$ and (some factor) of the target potential $V_{\pi}$. Loosely speaking, the time horizon increases as $\alpha_{d}$ in Assumption 3 increases.
In the next theorem, we show that the preservation of log-concavity under the FR flow can be exploited to ensure the same for the WFR flow uniformly in time, even when the W flow may not preserve log-concavity uniformly. The proof can be found in Appendix C.2.
@@
WFR preserves log-concavity uniformly in time. Assume the conditions of Lemma 4.1. Furthermore, if $\alpha_{h}-L_{\pi}<0$, assume that $b:=\sqrt{\frac{|\alpha_{h}-L_{\pi}|}{2}}$ satisfies
$b^{2}<\tfrac{\alpha_{\pi}}{4}.$ (4.4)
@@
$\alpha_{t}=\frac{\alpha_{\pi}}{2}-\frac{1}{4}+m\left(\frac{m_{0}-e^{-4mt}}{m_{0}+e^{-4mt}}\right),$ (4.5)
$m_{0}:=\frac{m+c_{0}+\tfrac{1}{4}}{m-c_{0}-\tfrac{1}{4}}$, $\enskip m=\sqrt{\frac{1}{2}\left(\frac{1}{8}+r\right)}$ and
$r$ $=\begin{cases}\frac{\alpha_{\pi}}{2}-2b^{2},\quad\alpha_{h}-L_{\pi}<0\\ \frac{\alpha_{\pi}}{2}+2b^{2},\quad\alpha_{h}-L_{\pi}>0\end{cases}$
and $c_{0}$ is given by (4.2). Furthermore, $\alpha_{\infty}>\frac{\alpha_{\pi}}{2}$.
The tightness of the constant (4.5) is evaluated empirically for 1D targets in Figure 4.1 where the true constants are known. In the Gaussian case, the constants are tight, so long as the assumptions are met. We also consider a non-Gaussian target, corresponding to a perturbation of a Gaussian $\pi(x)\propto\exp\left(-\frac{x^{2}}{10}-0.4\log(1+e^{x})\right)$. As can be seen there, (4.5) is reasonably tight, although its quality degrades as $t\rightarrow\infty$. Although $\alpha_{\infty}$ can be different from $\alpha_{\pi}$, the true asymptotic constant, it is never worse than $\frac{\alpha_{\pi}}{2}$.
Similar techniques can be used to obtain a uniform in $n$ preservation of log-concavity for the W-FR split scheme, so long as $\gamma$ is not too large, as shown in the following corollary. The proof is essentially identical to that in Appendix C.2, up until the limit $\tau_{i}\rightarrow 0$ is taken.
Corollary 4.2.
Assume the conditions of Theorem 4.1. Then $\nu_{n}$, the solution of the W-FR split scheme (2.6) with step size $\gamma$, is $\beta_{n}$-strongly log-concave uniformly in $n$ whenever $\alpha_{h}-L_{\pi}>0$. When $\alpha_{h}-L_{\pi}<0$, the same is true, as long as
$0<\gamma<\min(\tau_{1}^{\ast},\tau_{\infty}^{\ast})$
where $\tau_{1}^{\ast}$ and $\tau_{\infty}^{\ast}$ are as defined in (3.20). Furthermore, $\beta_{n}>\frac{\alpha_{\pi}}{2},\quad n=1,2,3,\dots$
We will rely on the results of Theorem 4.1 and Corollary 4.2 in the next section to characterise convergence rates.
Figure 4.1: Comparison of log-concavity constant from Theorem 4.1 and true log-concavity constant for the solution of (2.1) for 1D targets. Target densities considered are $\pi(x)=\mathcal{N}(x;,0,100)$ (left), $\pi(x)=\mathcal{N}(x;,0,2)$ (middle) and $\pi(x)\propto\exp\left(-\frac{x^{2}}{10}-0.4\log(1+e^{x})\right)$ (right). For the Gaussian targets, (3.1) is used to obtain the true $\alpha_{t}$ whereas for the non-Gaussian target, $\alpha_{t}$ is evaluated numerically from an Euler discretisation of (2.1). In all cases, $\mu_{0}(x)=\mathcal{N}(x;,0,1)$.
5 Convergence speed: the log-concave case
The preservation of log-concavity under the WFR flow and under the W-FR splitting established in the previous section can be exploited to obtain sharp rates of convergence to $\pi$ for both the exact WFR flow and the W-FR splitting. Existing $\mathrm{KL}$ decay results for the W flow and FR flow developed in the literature cannot be straightforwardly utilised to characterise differences in convergence speed between the W-FR and FR-W scheme. In particular, combining existing exponential decay bounds for the KL divergence for the W part and the FR part would not allow us to isolate changes to convergence speed depending on the order as the exponents simply add. Hence, the first variation results developed in Proposition 2.2 and 2.3 will be critical in characterising how convergence speed is affected due to splitting, and the ordering of operators. Due to the intractability of (2.12), we primarily focus on the W-FR scheme as a first step towards understanding the more general log-concave case. We first consider the exact WFR flow.
5.1 Exact WFR
The rate of decay of $\mathrm{KL}$ for the WFR flow is obtained considering the time derivative of $\mathrm{KL}(\mu_{t}||\pi)$ along the flow via classical arguments valid under log-concavity assumptions,
@@ 5.1 Exact WFR
As observed in [GM17, page 13], the first term corresponds to the negative gradient of $\mathrm{KL}$ w.r.t. the Wasserstein-2 metric, while the remaining terms give the negative gradient of $\mathrm{KL}$ w.r.t. the Fisher–Rao geometry, implying that the dissipation for the WFR metric is the sum of the W and the FR dissipation. As a consequence, the convergence rate of WFR gradient flow (2.1) can be upper bounded by $\mathrm{KL}(\mu_{t}||\pi)\leq\min\left\{\mathrm{KL}(\mu_{t}^{\textrm{FR}}|\pi),\mathrm{KL}(\mu_{t}^{\textrm{W}}|\pi)\right\},$ where $\mu_{t}^{\textrm{FR}},\mu_{t}^{\textrm{W}}$ denote the solutions of the FR and W flow PDEs respectively. A bound on the decay of KL along the W flow can be obtained if $\pi$ satifies an LSI (2.8) (e.g., [CEL+24]). Decay of KL along the FR flow at rate $\mathcal{O}(e^{-t})$ is guaranteed under boundedness conditions on $\log\pi/\mu_{0}$ ([CHH+23, Eq. (B.6)]), leading to
$\mathrm{KL}(\mu_{t}||\pi)\leq\min\left\{e^{-2\lambda_{\pi}^{-1}t}\mathrm{KL}(\mu_{0}||\pi),Me^{-t}(2+B+Be^{Me^{-t}(1+B)})\right\},$ (5.2)
with $\int|x|^{2}\mu_{0}(x)dx\leq B$, $\int|x|^{2}\pi(x)dx\leq B$ and $\left|\log\nicefrac{{\mu_{0}(x)}}{{\pi(x)}}\right|\leq M(1+|x|^{2})$.
A sharper rate of converge for WFR can be obtained assuming $\log\nicefrac{{\pi(x)}}{{\mu_{0}(x)}}\geq M$ and a warm-start condition [LLN19, Appendix B]
@@
for all $t\geq t_{0}:=\log(M/\delta^{3})$ and $\delta>0$. We point out that neither (5.2) nor (5.3) provide a sharp rate of decay for $\mathrm{KL}(\mu_{t}||\pi)$. In fact, (5.2) shows that the rate of WFR is never worse than that of W or FR, looking at (5.1), however it is evident that the rate of decay of the WFR flow should be considerably higher than that of the two singular flows. Similarly, (5.3) shows that the decay in the case of the WFR flow is faster than that of the W flow but the rate depends on an arbitrary constant $\delta$ which makes (5.3) harder to interpret.
In Proposition 5.1, we obtain a result that shows the decay rate is indeed the sum of the W and FR rates, but for the symmetrised KL, not (reverse) KL. The symmetrised $\mathrm{KL}$ divergence, or Jeffrey’s divergence $J$, is defined as
@@
for any $\mu\ll\pi$ and also $\pi\ll\mu$. The rationale behind considering the symmetrised KL is due to the fact that the KL is not geodesically convex under the FR flow [CCH+24, Theorem 1.1] nor satisfies a gradient dominance condition [CCH+24, Theorem 4.1], which makes its convergence analysis more difficult. On the other hand, Jeffrey’s divergence satisfies a gradient dominance condition [CCH+24, Section 4] which allows to achieve improved rates. In addition, $J(\mu,\pi)$ upper bounds $\mathrm{KL}(\mu||\pi)$ and thus the obtained rates for $J(\mu,\pi)$ gives information on the decay of KL too. The decay rate obtained in Proposition 5.1 is a sum of the decay rates of W and FR flows111where the constant for the W flow is instead written in terms of the LSI constant, since W flows preserve LSI uniformly in time [LMW25, Lemma 16]., which was conjectured by [DP23], but not yet proven in the literature.
Theorem 4.1 immediately provides sufficient conditions to establish an exponential decay result of the symmetrised $\mathrm{KL}$ divergence, or Jeffrey’s divergence $J$. The proof makes use of recent functional inequality results on pure FR flows [CCH+24, Section 4] and provides a sharper rate of convergence than the results available in the literature (see Figure 5.1 for a Gaussian example).
@@
Assume the conditions of Theorem 4.1. Then the following decay result holds for $\mu_{t}$ the solution of the WFR PDE (2.1) for all $t>0$
$J(\mu_{t},\pi)\leq J(\mu_{0},\pi)e^{-t(\alpha_{\pi}+1)}.$
@@
Given the assumptions, we may interchange differentiation and integration to obtain
@@
$\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)(x)f_{\textrm{W}}(\mu_{t}(x))dx$ $=\int\left(\log\frac{\mu_{t}(x)}{\pi(x)}-\frac{\pi(x)}{\mu_{t}(x)}\right)\nabla\cdot\left(\mu_{t}(x)\nabla\log\frac{\mu_{t}}{\pi}(x)\right)(x)dx$
@@
$\leq-2\min(\lambda_{\pi}^{-1},\lambda_{\mu_{t}}^{-1})(\mathrm{KL}(\mu_{t}||\pi)+\mathrm{KL}(\pi||\mu_{t}))$ (5.5)
@@
$<-\alpha_{\pi}J(\mu_{t},\pi),$ (5.6)
where $\mathcal{I}(\mu_{t}||\pi)$ is the relative Fisher information between $\mu_{t}$ and $\pi$. The inequality (5.5) holds due to Assumption 2 and Theorem 4.1. More specifically, since $\pi$ is $\alpha_{\pi}$-strongly log-concave, it immediately satisfies a Log-Sobolev inequality (2.8) with constant $\lambda_{\pi}=\alpha_{\pi}^{-1}$ (likewise for $\mu_{t}$, which is $\alpha_{t}$-strongly log-concave due to Theorem 4.1). The final inequality (5.6) holds since due to Theorem 4.1, $\alpha_{t}$ is monotonic in $t$ and $\alpha_{\infty}>\frac{\alpha_{\pi}}{2}$ and finally under Assumption 3(a), $\alpha_{0}>\frac{\alpha_{\pi}}{2}$ so that $\alpha_{t}>\frac{\alpha_{\pi}}{2}$ for all $t\geq 0$.
Then for the second term in (5.4), by a direct application of (5.16) in [CCH+24, Theorem 5.6] with $f(y)=y\log(y)$ and $\bar{f}=yf(y^{-1})$, it holds that
$-\int\frac{\pi(x)}{\mu_{t}(x)}f_{\textrm{FR}}(\mu_{t}(x))dx=-J(\mu_{t},\pi),$
@@
$\int\log\frac{\mu_{t}(x)}{\pi(x)}f_{\textrm{FR}}(\mu_{t}(x))dx=-\text{Var}_{\mu_{t}}\left(\log\frac{\mu_{t}}{\pi}\right)<0,$
@@
$\int\left(\log\frac{\mu_{t}}{\pi}-\frac{\pi}{\mu_{t}}\right)(x)f_{\textrm{FR}}(\mu_{t}(x))dx<-J(\mu_{t},\pi).$ (5.7)